# Mechanistic Interpretability for AI Safety A Review

### v1

@@ 3.2 Computation and Abstraction
World models are internal causal models of an environment formed within neural networks. Traditionally linked with reinforcement learning, these models are explicitly trained to develop a compressed spatial and temporal representation of the training environment, enhancing downstream task performance and sample efficiency through training on internal hallucinations (Ha & Schmidhuber, 2018). However, in the context of our survey, our focus shifts to world models that potentially form implicitly as a by-product of the training process, especially in LLMs that are trained on next-token prediction - also called GPT.
@@
A critical perspective often surfacing in discussions about LLMs is their characterization as stochastic parrots (Bender et al., 2021). This label stems from their fundamental operational mechanism of predicting the next word in a sequence, supposedly relying heavily on memorization. From this viewpoint, LLMs are seen as forming complex correlations based on observational data but are thought to lack the ability to develop causal models of the world. This limitation is attributed to their lack of access to interventional data (Pearl, 2009).
@@
However, this understanding of LLMs shifts significantly when viewed through the lens of the active inference framework (Salvatori et al., 2023), a theory rooted in cognitive science and neuroscience. Active inference postulates that the objective of minimizing prediction error, given enough representative capacity, is adequate for a learning system to develop complex world representations, behaviors, and abstractions. Since language inherently mirrors the world, these models could implicitly construct linguistic and broader world models. This perspective presupposes that LLMs, in their pursuit of better modeling sequences, inherently learn world models, abstractions, and algorithms for this purpose (Kulveit et al., 2023).
@@
This alternative understanding of LLMs aligns with the simulation hypothesis, which suggests that models designed for prediction, such as LLMs, will eventually simulate the causal processes underlying data creation. Seen as an extension of their drive for efficient compression, this hypothesis implies that adequately trained models like GPT could develop internal world models as a natural outcome of their predictive training (janus, 2022).
@@
[htb] Hypothesis: Simulation A model whose objective is text prediction will simulate the causal processes underlying the text creation if optimized sufficiently strongly (janus, 2022).
@@
In addition to theoretical considerations for emergent causal world models (Richens & Everitt, 2024; Nichani et al., 2024), mechanistic interpretability is starting to provide empirical evidence on the types of internal world models that may emerge in LLMs. The ability to internally represent the board state in games like Othello (Li et al., 2023a; Nanda et al., 2023b), create linear abstractions of spatial and temporal data (Gurnee & Tegmark, 2023), and structure complex representations of mazes, demonstrating an understanding of maze topology and pathways (Ivanitskiy et al., 2023) highlight the growing abstraction capabilities of LLMs.
These emergent world models have significant implications for AI alignment research. For example, finding an internal representation of human values and aiming the AI systems objective may be the most trivial way to achieve alignment (Wentworth, 2022). Especially if the world model is internally separated from notions of goals and agency (Ruthenis, 2022), world model interpretability may be enough for alignment (Ruthenis, 2023).
Conditioning of pre-trained models as a pathway towards general intelligence is deemed comparatively safe, as it avoids directly creating agents with inherent goals or agendas (Jozdien, 2022; Hubinger et al., 2023). However, Hubinger et al. (2023) also highlights that prompting a model to simulate an actual agent, such as "You are a superintelligence in 2035 writing down an alignment solution:", could inadvertently lead to the formation of internal agents. In contrast, training with reinforcement learning tends to create agents by default (Casper et al., 2023a; Ngo et al., 2022).
@@
The prediction orthogonality hypothesis further expands on this idea: It posits that prediction-focused models like GPT may simulate agents with various objectives and levels of optimality. In this context, GPT are simulators, simulating entities known as simulacra that can be either agentic or non-agentic, with different objectives from the simulator itself (janus, 2022).
@@
[htb] Hypothesis: Prediction Orthogonality A model whose objective is prediction can simulate agents who optimize toward any objectives with any degree of optimality (janus, 2022).
This prediction orthogonality hypothesis suggests that models primarily focused on prediction, such as GPT, can simulate agents — referred to as ’simulacra’ — with potentially misaligned objectives (janus, 2022). Although GPT may lack genuine agency or intentionality, it may produce outputs that simulate these qualities (Bereska & Gavves, 2023), underscoring the need for careful oversight and, better yet, finding internal agents or their constituents such as optimization or search potentially via mechanistic interpretability - an endeavor also known as searching for search (NicholasKees & janus, 2022).
@@ 4.1 Observation
Structured Probes
While most of this review focuses on bottom-up, mechanistic approaches to interpretability, it is worth considering the potential for integrating top-down, concept-based techniques like structured probes. Structured probes represent an advanced technique in conceptual interpretability, playing a crucial role in probing language models to uncover complex features like truth representations.
A notable advancement in this domain is the discovery of an "internal truth" direction within language models using unsupervised contrastive probing, as demonstrated by the contrast-consistent search (CCS) method proposed by Burns et al. (2023). CCS identifies linear projections of hidden states that exhibit logical consistency, ensuring contrasting truth values for statements and their negations, contributing substantially to conceptual interpretability (Zou et al., 2023).
However, structured probes face significant challenges, particularly in unsupervised probing scenarios. A major concern is verifying the accuracy of discovered features, as unsupervised methods can identify numerous features without a straightforward verification process. Additionally, recent work by Farquhar et al. (2023) raises doubts about the scalability of CCS, suggesting that the CCS loss may capture simulated knowledge rather than the model’s true knowledge, especially in highly capable models adept at simulating agents (simulacra) with their own belief systems.
While structured probes primarily focus on high-level conceptual representations, their findings could potentially inform or complement mechanistic interpretability efforts. For instance, identifying truth directions through structured probes could help guide targeted interventions or analyze the underlying circuits responsible for truthful behavior using mechanistic techniques like activation patching or circuit tracing (Section 4.2). Conversely, mechanistic methods could provide insights into how truth representations emerge and are computed within the model, addressing some of the challenges faced by unsupervised structured probes.
Logit Lens
@@
Feature Disentanglement via Sparse Dictionary Learning
@@
Practical implementations, such as the toy model by Sharkey et al. (2022b), demonstrate the viability of this approach, with the precise tuning of the sparsity penalty on the hidden activations being a critical aspect that dictates the sparsity level of the autoencoder (Sharkey et al., 2022b). We show an overview in the pink box on sparse autoencoders in Figure 6.
[htb] Sparse Dictionary Learning Sparse autoencoders (Cunningham et al., 2024) are a solution to the sparse dictionary learning (Olshausen & Field, 1997) problem to decompose neural network activations into individual component features. The goal is to learn a dictionary of vectors $\{\mathbf{f}_{k}\}_{k=1}^{n_{\text{feat}}}\subset\mathbb{R}^{d}$ that can represent the unknown, ground truth network features $\{\mathbf{g}_{j}\}_{j=1}^{n_{\text{gt}}}$ as sparse linear combinations. The autoencoder architecture consists of an encoder and a ReLU activation function, expanding the input dimensionality to $d_{\text{hid}}=Rd_{\text{in}}$, where $R$ controls the ratio of the feature dictionary size to the model dimension. The encoder’s output is given by: $\mathbf{h}$ $=\text{ReLU}(W_{\text{enc}}\mathbf{x}+\mathbf{b})$ (1) ${\mathbf{x^{\prime}}}$ $=W_{\text{dec}}\mathbf{h}=\sum_{i=0}^{d_{\text{hid}}-1}h_{i}\mathbf{f}i$ (2) where $W_{\text{enc}},W_{\text{dec}}^{T}\in\mathbb{R}^{d_{\text{hid}}\times d_{\text{in}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{\text{hid}}}$. The parameter matrix $W_{\text{dec}}$ forms the feature dictionary, with rows $\mathbf{f}_{i}$ as dictionary features. The autoencoder is trained to minimize the loss, where the ${L}_{1}$ penalty on $\mathbf{h}$ encourages sparse reconstructions using the dictionary features. $\mathcal{L}(\mathbf{x})=||\mathbf{x}-{\mathbf{x^{\prime}}}||_{2}^{2}+\alpha||\mathbf{h}||_{1}$ (3) [Uncaptioned image] Figure 6: Illustration of a sparse autoencoder applied to the MLP layer activations, consisting of an encoder that increases dimensionality while emphasizing sparse representations and a decoder that reconstructs the original activations using the learned feature dictionary.
Empirical studies indicate that sparse autoencoders can enhance the interpretability of neural networks, exhibiting higher scores on the autointerpretability metric and increased monosemanticity (Bricken et al., 2023; Cunningham et al., 2024; Sharkey et al., 2022b). Furthermore, sparse autoencoders have been employed to measure feature sparsity (Deng et al., 2023) and interpret reward models in reinforcement learning-based language models (Marks et al., 2023), making them an actively researched area in mechanistic interpretability.
@@ 4.2 Intervention
Activation Patching
is a collective term for a set of causal intervention techniques, also known as causal tracing (Meng et al., 2022a), interchange intervention (Geiger et al., 2021b), causal mediation analysis (Vig et al., 2020), and causal ablation (Wang et al., 2023). While nuanced in their application, these techniques share the common goal of manipulating neural network activations to shed light on the decision-making processes within the model.
Activation patching modifies a neural model’s internal state by replacing specific activations with alternative values, such as zeros, mean activations across samples, random noise, or activations from a different forward pass. This technique enables researchers to isolate and examine the effects of modifying particular neural circuits within the model to understand how these circuits interact and contribute to the model’s behavior in response to different inputs. By selectively replacing activations, activation patching highlights the circuits directly responsible for particular behaviors or outputs while reducing the influence of other, irrelevant circuits.
The primary goal is to isolate and understand the role of specific components or circuits within the model. By observing how changes in activations affect the model’s output, researchers can infer the function and importance of those components. Critical applications are (i) localizing behavior by identifying critical activations, for example, understanding storage and processing of factual information (Meng et al., 2022a; Geva et al., 2023; Goldowsky-Dill et al., 2023; Stolfo et al., 2023), and (ii) analyzing component interactions, such as conducting circuit analysis to identify sub-networks within a model’s computation graph that implement specified behaviors (Wang et al., 2023; Hanna et al., 2023; Lieberum et al., 2023; Hendel et al., 2023; Geva et al., 2023).
(a) The activation patching process
(b) Boolean logic analogies
Figure 7: (a) The transfer of activations from clean to corrupted inputs isolates neural circuits. (b) Boolean logic circuits are an analogy for sufficiency and necessity in neural circuits via patching strategies.
The standard protocol of activation patching entails: i.) running a model with a clean input and caching the latent activations; ii.) executing the model with a corrupted input; iii.) re-running the model with the corrupted input but substituting specific activations with those from the clean cache; and iv.) determining significance by observing the variations in the model’s output during the third step, thereby highlighting the importance of the replaced components.
The process relies on comparing pairs of inputs: a clean input, which triggers the desired behavior, and a corrupted input, which is identical to the clean one except for critical differences that prevent the behavior. This careful selection ensures that the two inputs share as many circuits as possible, except those directly influencing the behavior under investigation - effectively controlling for confounding circuitry. Through activation patching—transferring activations from the clean input run to the corrupted one—researchers can maintain the shared circuits’ functionality while pinpointing and isolating the specific circuit responsible for the behavior.
Differences in patching direction — clean to corrupted versus corrupted to clean — yield insights into which model components are sufficient or necessary for a given behavior. Clean to corrupted patching (causal tracing) identifies activations that are sufficient for restoring clean performance, highlighting the non-necessity of redundant components in achieving specific outputs. If many components redundantly encode something that quickly saturates, you can get good performance from patching in any of them, even if none are necessary. This approach, effective even in redundant system components, clarifies the sufficiency of specific activations in driving model performance under OR logic conditions: In circuits A-AND-B, this tells us nothing, but in A-OR-B, it tells us that both A or B is sufficient on its own.
Conversely, corrupted to clean patching (resample ablation) focuses on determining the necessary activations for clean performance, emphasizing the criticality of specific components. If the model has redundancy, we may see that nothing is necessary! Even if, in the aggregate, they’re essential. Particularly useful in AND logic scenarios, this method assesses the impact of removing specific activations, revealing the indispensable elements of the computational architecture. In the circuit A-OR-B, resample ablating does nothing. However, A-AND-B tells us that removing each of A or B will dramatically reduce performance.
Activation patching can employ various corruption methods, including zero-, mean-, random-, or resample ablation - replacing activations with zeros, an average over activations across diverse samples, Gaussian noise (Meng et al., 2022a), or activations from a different model run (Vig et al., 2020; Wang et al., 2023), each serving to modulate the model’s internal state in distinct ways. Among these, resample ablation stands out for its effectiveness in maintaining consistent model behavior by not changing the data distribution too much (Zhang & Nanda, 2023). While breaking behavior is always possible by taking the model off-distribution, this is uninteresting for finding the relevant circuit (Nanda, 2023e). Therefore, one needs to be careful when interpreting the results of patching.
Among evaluation metrics for assessing how activation patching influences behavior, comparing the logit difference between clean and corrupted runs stands out as a precise measure of the changes in confidence levels across different inputs and the ability to detect negative modules (Zhang & Nanda, 2023). Additionally, per-token log probability provides a detailed view of the model’s prediction confidence at each token, providing more granularity. Direct logit attribution further delves into how different components influence the logit of the correct next token, shedding light on the critical elements of the model’s predictions. However, internal memory management can mislead this metric (Dao et al., 2023). When used together, these metrics enable a thorough evaluation of the impact of activation patching, offering comprehensive insights into the intervention’s effects on model behavior.
Activation patching, while a powerful interpretability tool, encounters several limitations. A primary concern is its limited ability to generalize beyond specific distributions or tasks, often focusing on narrow scenarios without fully addressing broader or varied contexts (Nanda, 2023d). Another issue is the MLP-In-The-Middle illusion (Lange et al., 2023), a phenomenon where patching an entire Multi-Layer Perceptron (MLP) layer shows no observable effect, yet patching a specific subspace within the same layer reveals significant impacts. This raises questions about the relevance of certain subspaces in the model’s normal functioning. This suggests that some components may appear crucial in patching but are dormant or irrelevant in regular operations.
Additionally, the Hydra effect (McGrath et al., 2023), where models internally self-repair and maintain capabilities even when key components are ablated, can sometimes obscure the relevant components. The effects of patching can propagate and interact in complex ways, potentially exaggerating or diminishing the apparent importance of certain components (see Section 8.2).
Translating the localization of model behaviors, as revealed by activation patching, into effective model editing (Hase et al., 2023) can also be challenging. Understanding where certain information or processes are stored in the model doesn’t always seem to translate into actionable strategies for modifying or improving the model’s performance or behavior.
Furthermore, the process of activation patching can be slow, which is especially problematic in large models or when attempting to automate the process (Conmy et al., 2023). However, this challenge can be partially mitigated using attribution patching (Nanda, 2023d; Syed et al., 2023), a gradient-based alternative that takes a linear approximation to traditional activation patching, similar to other attribution methods (see Section 2). Attribution patching offers a faster and more scalable approach, particularly advantageous in automated circuit discovery (Syed et al., 2023) and large model applications, providing a more feasible and efficient means of probing neural network behaviors while retaining the core benefits of the activation patching approach.
Recent advancements include the introduction of AtP* (Kramár et al., 2024), a refined version of attribution patching that addresses specific failure modes of the original method to reduce false negatives, thus improving its reliability while maintaining scalability. Other variations include path patching (Goldowsky-Dill et al., 2023), which quantitatively tests hypotheses expressing that behaviors are localized to a set of paths, and attention pattern patching (Nanda, 2023d), which leverages attention attribution patterns to gain insights into information flow within the network. Ghandeharioun et al. (2024) introduced a unified framework to analyze hidden representations.
4.3 Integrating Observation and Intervention
Integrating different methodologies is necessary for a thorough understanding of neural network models. These complex models require a broad approach that goes beyond individual techniques. An effective strategy combines feature-level analysis tools like sparse autoencoders with probing and interventional methods like activation patching. This integrated approach could allow a more in-depth examination of neural network feature-level circuits.
Even seemingly independent techniques can improve the robustness of analysis. For example, validating activation patching findings with maximally activating dataset examples or direct logit attribution offers a more comprehensive view of a component’s network functionality. However, achieving complete understanding remains challenging due to the potential for feature superposition within these models. A single component may simultaneously represent multiple features, complicating interpretability efforts. Navigating and disentangling these intertwined representations requires integrating diverse analytical techniques.
5 Evaluation
Qualitative Evaluation.
Interpretability research lacks established metrics, making qualitative results crucial. The signal of structure approach (Olah & Jermyn, 2024) – observing intricate patterns indicating genuine structures – resembles examining cells under a microscope. A nuanced balance of qualitative observations and quantitative analyses is required, often necessitating custom interfaces to avoid oversimplification.
Quantitative Evaluation.
A central challenge is the lack of rigorous evaluation methods. Relying solely on intuition is inadequate, as hypotheses can be conflated with conclusions (Rudin, 2019; Miller, 2019; Räuker et al., 2023), leading to cherry-picking and optimizing for best-case performance rather than aiming for methods that perform well on average or in worst-case scenarios (Casper, 2023) (see also Section 8.1). Current practices are ad hoc, with proxies (Doshi-Velez & Kim, 2017) potentially leading to over-optimization (Goodhart’s law - When a measure becomes a target, it ceases to be a good measure). Distinguishing correlation from causation is crucial, as interpretability illusions (Bolukbasi et al., 2021; Friedman et al., 2023a; Olah et al., 2017) demonstrate visualizations may be meaningless without causal linking.
Rigorous evaluation methods are needed, such as i.) assessing out-of-distribution inputs, as most current methods are only valid for analyzing specific examples or datasets (Räuker et al., 2023; Ilyas et al., 2019; Mu & Andreas, 2020; Casper et al., 2023b; Burns et al., 2023), ii.) controlling the system through edits, such as implanting or removing trojans (Mazeika et al., 2022) or targeted editing (Ghorbani & Zou, 2020; Dai et al., 2022; Meng et al., 2022a; b; Bau et al., 2018; Hase et al., 2023), iii.) or replacing it with simpler reverse-engineered alternatives (Lindner et al., 2023). Ultimately, establishing benchmarks, ideally automated, is required.
@@ 5 Evaluation
The theory of causality (Pearl, 2009) provides a mathematically precise language for mechanistic interpretability, forming the foundation for understanding high-level semantics in neural representations (Geiger et al., 2023a). Treating neural networks as causal models involves considering the compute graph as the causal graph, allowing for precise interventions and examining individual parameters’ roles (McGrath et al., 2023). In contrast to typical real-world causal analyses, the causal model is known with complete certainty, long chains of interventions are possible, and all variable values can be simultaneously read. Nevertheless, the large number of parameters, often lacking clear meaning, poses a challenge in this context.
Causal inference techniques have been employed in various contexts within neural networks, including locating factual knowledge (Meng et al., 2022a), addressing gender bias through mediation analyses (Vig et al., 2020), and constructing causal abstractions of neural network computations (Geiger et al., 2023a; 2021a; b; 2021b; McGrath et al., 2023). Ablations and interchange interventions have been suggested as means to validate hypotheses about mechanisms in neural networks and enforce specific structures (Chan et al., 2022; Leavitt & Morcos, 2020; Geiger et al., 2021b), enabling large-scale analysis of model behavior (Wu et al., 2023).
Rigorous Hypothesis Testing.
Causal scrubbing (Chan et al., 2022), causal abstraction (Geiger et al., 2023a), and locally consistent abstractions (Jenner et al., 2023) have been proposed as rigorous methods to formalize and test hypotheses about how neural networks implement specific behaviors.

### v2

@@ 3.4 Emergence of World Models and Simulated Agents
World models are internal causal models of an environment formed within neural networks. Traditionally linked with reinforcement learning, these models are explicitly trained to develop a compressed spatial and temporal representation of the training environment, enhancing downstream task performance and sample efficiency through training on internal hallucinations (Ha & Schmidhuber, 2018). However, in the context of our survey, our focus shifts to internal world models that potentially form implicitly as a by-product of the training process, especially in LLMs trained on next-token prediction – also called GPT.
@@
LLMs are sometimes characterized as stochastic parrots (Bender et al., 2021). This label stems from their fundamental operational mechanism of predicting the next word in a sequence, which is seen as relying heavily on memorization. From this viewpoint, LLMs are thought to form complex correlations based on observational data but lack the ability to develop causal models of the world due to their lack of access to interventional data (Pearl, 2009).
@@
An alternative perspective on LLMs comes from the active inference framework (Salvatori et al., 2023), a theory rooted in cognitive science and neuroscience. Active inference postulates that the objective of minimizing prediction error, given enough representative capacity, is adequate for a learning system to develop complex world representations, behaviors, and abstractions. Since language inherently mirrors the world, these models could implicitly construct linguistic and broader world models.
@@
The simulation hypothesis suggests that models designed for prediction, such as LLMs, will eventually simulate the causal processes underlying data creation. Seen as an extension of their drive for efficient compression, this hypothesis implies that adequately trained models like GPT could develop internal world models as a natural outcome of their predictive training (janus, 2022).
@@
[htb] Hypothesis 5: Simulation A model whose objective is text prediction will simulate the causal processes underlying the text creation if optimized sufficiently strongly (janus, 2022).
@@
In addition to theoretical considerations for emergent causal world models (Richens & Everitt, 2024; Nichani et al., 2024), mechanistic interpretability is starting to provide empirical evidence on the types of internal world models that may emerge in LLMs. The ability to internally represent the board state in games like Othello (Li et al., 2023a; Nanda et al., 2023b), create linear abstractions of spatial and temporal data (Gurnee & Tegmark, 2024), and structure complex representations of mazes, demonstrating an understanding of maze topology and pathways (Ivanitskiy et al., 2023) highlight the growing abstraction capabilities of LLMs. Li et al. (2021) identified contextual word representations that function as models of entities and situations evolving throughout a discourse, akin to linguistic models of dynamic semantics. Patel & Pavlick (2022) demonstrated that LLMs can map conceptual domains (e.g., direction, color) to grounded world representations given a few examples, suggesting they learn rich conceptual spaces (Gardenfors, 2004) reflective of the non-linguistic world.
@@
The prediction orthogonality hypothesis further expands on this idea: It posits that prediction-focused models like GPT may simulate agents with various objectives and levels of optimality. In this context, GPT are simulators, simulating entities known as simulacra that can be either agentic or non-agentic, with different objectives from the simulator itself (janus, 2022). The implications of the simulation and prediction orthogonality hypotheses for AI safety and alignment are discussed in Section 6.
@@
[htb] Hypothesis 6: Prediction Orthogonality A model whose objective is prediction can simulate agents who optimize toward any objectives with any degree of optimality (janus, 2022).
@@ 4.1 Observation
Structured Probes.
While focusing on bottom-up, mechanistic interpretability approaches, we can also consider integrating top-down, concept-based structured probes with mechanistic interpretability.
Structured probes aid conceptual interpretability, probing language models for complex features like truth representations. Notably, Burns et al. (2023)’s contrast-consistent search (CCS) method identifies linear projections exhibiting logical consistency in hidden states, contrasting truth values for statements and negations.
However, structured probes face significant challenges in unsupervised probing scenarios. As Farquhar et al. (2023) showed, arbitrary features, not just knowledge-related ones, can satisfy the CCS loss equally well, raising doubts about scalability. For example, the loss may capture simulation of knowledge from hypothesized simulacra within sufficiently powerful language models rather than the models’ true knowledge. Furthermore, Farquhar et al. (2023) demonstrates unsupervised methods like CCS often detect prominent but unintended features in the data, such as distractors like "banana." The discovered features are also highly sensitive to prompt choice, and there is no principled way to select prompts that would reliably surface a model’s true knowledge.
While structured probes primarily focus on high-level conceptual representations (Zou et al., 2023), their findings could potentially inform or complement mechanistic interpretability efforts. For instance, identifying truth directions through structured probes could help guide targeted interventions or analyze the underlying circuits responsible for truthful behavior using mechanistic techniques such as activation patching or circuit tracing (Section 4.2). Conversely, mechanistic methods could provide insights into how truth representations emerge and are computed within the model, addressing some of the challenges faced by unsupervised structured probes.
Logit Lens.
@@
[htb] Sparse Dictionary Learning Sparse autoencoders (Cunningham et al., 2024) represent a solution attempt to the challenge of polysemantic neurons. The problem of superposition is mathematically formalized as sparse dictionary learning (Olshausen & Field, 1997) problem to decompose neural network activations into disentangled component features. The goal is to learn a dictionary of vectors $\{\mathbf{f}_{k}\}_{k=1}^{n_{\text{feat}}}\subset\mathbb{R}^{d}$ that can represent the unknown, ground truth network features as sparse linear combinations. If successful, the learned dictionary contains monosemantic neurons corresponding to features (Sharkey et al., 2022b). The autoencoder architecture consists of an encoder and a ReLU activation function, expanding the input dimensionality to $d_{\text{hid}}>d_{\text{in}}$. The encoder’s output is given by: $\mathbf{h}$ $=\text{ReLU}(W_{\text{enc}}\mathbf{x}+\mathbf{b})$ (1) ${\mathbf{x^{\prime}}}$ $=W_{\text{dec}}\mathbf{h}=\sum_{i=0}^{d_{\text{hid}}-1}h_{i}\mathbf{f}i$ (2) where $W_{\text{enc}},W_{\text{dec}}^{T}\in\mathbb{R}^{d_{\text{hid}}\times d_{\text{in}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{\text{hid}}}$. The parameter matrix $W_{\text{dec}}$ forms the feature dictionary, with rows $\mathbf{f}_{i}$ as dictionary features. The autoencoder is trained to minimize the loss, where the ${L}_{1}$ penalty on $\mathbf{h}$ encourages sparse reconstructions using the dictionary features. $\mathcal{L}(\mathbf{x})=||\mathbf{x}-{\mathbf{x^{\prime}}}||_{2}^{2}+\alpha||\mathbf{h}||_{1}$ (3) [Uncaptioned image] Figure 8: Illustration of a sparse autoencoder applied to the MLP layer activations, consisting of an encoder that increases dimensionality while emphasizing sparse representations and a decoder that reconstructs the original activations using the learned feature dictionary.
Feature Disentanglement via Sparse Dictionary Learning.
@@
Practical implementations, such as the toy model by Sharkey et al. (2022b), demonstrate the viability of this approach, with the precise tuning of the sparsity penalty on the hidden activations being a critical aspect that dictates the sparsity level of the autoencoder. We show an overview in the pink box on sparse autoencoders in Figure 8.
Empirical studies indicate that sparse autoencoders can enhance the interpretability of neural networks, exhibiting higher scores on the autointerpretability metric and increased monosemanticity (Bricken et al., 2023; Cunningham et al., 2024; Sharkey et al., 2022b). Furthermore, sparse autoencoders have been employed to measure feature sparsity (Deng et al., 2023) and interpret reward models in reinforcement learning-based language models (Marks et al., 2023a), making them an actively researched area in mechanistic interpretability.
@@ 4.2 Intervention
The theory of causality (Pearl, 2009) provides a mathematically precise framework for mechanistic interpretability, offering a rigorous approach to understanding high-level semantics in neural representations (Geiger et al., 2023a). By treating neural networks as causal models, with their compute graphs serving as causal graphs, researchers can perform precise interventions and examine the roles of individual parameters (McGrath et al., 2023). This causal perspective on interpretability has led to the development of various intervention techniques, including activation patching (Section 4.2.1), causal abstraction (Section 4.2.2), and hypothesis testing methods (Section 4.2.3).
4.2.1 Activation Patching
Figure 9: (a) Activation patching in a transformer model. Left: The model processes the clean input "Colosseum in Rome," caching the latent activations (step i). Right: The model runs with the corrupted input "Eiffel Tower in Paris" (step ii). The pink arrow shows an MLP layer activation (green diamond) patched from the clean run into the corrupted run (step iii). This causes the prediction to change from "Paris" to "Rome," demonstrating how the significance of the patched component is determined (step iv). By comparing these carefully selected inputs, researchers can control for confounding circuitry and isolate the specific circuit responsible for the location prediction behavior. (b) Activation patching directions: Top: Patching corrupted activations (orange) into clean circuits (turquoise) reveals sufficient components for identifying OR logic scenarios. Bottom: Patching clean activations (green) into corrupted circuits (orange) reveals necessary components that are useful for identifying AND logic scenarios. The AND and OR gates demonstrate how these patching directions uncover different logical relationships between model components.
Activation patching is a collective term for a set of causal intervention techniques that manipulate neural network activations to shed light on the decision-making processes within the model. These techniques, including causal tracing (Meng et al., 2022a), interchange intervention (Geiger et al., 2021b), causal mediation analysis (Vig et al., 2020), and causal ablation (Wang et al., 2023), share the common goal of modifying a neural model’s internal state by replacing specific activations with alternative values, such as zeros, mean activations across samples, random noise, or activations from a different forward pass (Figure 9a).
The primary objective of activation patching is to isolate and understand the role of specific components or circuits within the model by observing how changes in activations affect the model’s output. This enables researchers to infer the function and importance of those components. Key applications include localizing behavior by identifying critical activations, such as understanding the storage and processing of factual information (Meng et al., 2022a; Geva et al., 2023; Goldowsky-Dill et al., 2023; Stolfo et al., 2023), and analyzing component interactions through circuit analysis to identify sub-networks within a model’s computation graph that implement specified behaviors (Wang et al., 2023; Hanna et al., 2023; Lieberum et al., 2023; Hendel et al., 2023; Geva et al., 2023).
The standard protocol for activation patching (Figure 9a) involves: (step i) running the model with a clean input and caching the latent activations; (step ii) executing the model with a corrupted input; (step iii) re-running the model with the corrupted input but substituting specific activations with those from the clean cache; and (step iv) determining significance by observing the variations in the model’s output during the third step, thereby highlighting the importance of the replaced components. This process relies on comparing pairs of inputs: a clean input, which triggers the desired behavior, and a corrupted input, which is identical to the clean one except for critical differences that prevent the behavior. By carefully selecting these inputs, researchers can control for confounding circuitry and isolate the specific circuit responsible for the behavior.
Differences in patching direction—clean to corrupted (causal tracing) versus corrupted to clean (resample ablation)—provide insights into the sufficiency or necessity of model components for a given behavior. Clean to corrupted patching identifies activations sufficient for restoring clean performance, even if they are unnecessary due to redundancy, which is particularly informative in OR logic scenarios (Figure 9b, A-OR-B). Conversely, corrupted to clean patching determines the necessary activations for clean performance, which is useful in AND logic scenarios (Figure 9b, A-AND-B).
Activation patching can employ corruption methods, including zero-, mean-, random-, or resample ablation, each modulating the model’s internal state in distinct ways. Resample ablation stands out for its effectiveness in maintaining consistent model behavior by not changing the data distribution too much (Zhang & Nanda, 2023). However, it is essential to be careful when interpreting the patching results, as breaking behavior by taking the model off-distribution is uninteresting for finding the relevant circuit (Nanda, 2023e).
Path Patching and Subspace Activation Patching.
Path patching extends the activation patching approach to multiple edges in the computational graph (Wang et al., 2023; Goldowsky-Dill et al., 2023), allowing for a more fine-grained analysis of component interactions. For example, path patching can be used to estimate the direct and indirect effects of attention heads on the output logits. Subspace activation patching, also known as distributed interchange interventions (Geiger et al., 2023b), aims to intervene only on linear subspaces of the representation space where features are hypothesized to be encoded, providing a tool for more targeted interventions.
Limitations and Advancements.
Activation patching has several limitations, including the effort required to design input templates and counterfactual datasets, the need for human inspection to isolate important subgraphs, and potential second-order effects that can complicate the interpretation of results (Lange et al., 2023) and the hydra effect (McGrath et al., 2023) (see discussion in Section 7.2). Recent advancements aim to address these limitations, such as automated circuit discovery algorithms (Conmy et al., 2023), gradient-based methods for scalable component importance estimation like attribution patching (Nanda, 2023d; Syed et al., 2023), and techniques to mitigate self-repair interferences during analysis (Ferrando & Voita, 2024).
4.2.2 Causal Abstraction